# F&O Expiry vs. First-Day SIPs: A 22-Year Analysis of Timing Advantages in India’s Nifty 50

### v1

@@ Abstract
This study presents a comparative analysis of Systematic Investment Plan (SIP) performance in India’s Nifty 50 index over a 22-year period (2003–2024), focusing on the timing of investments—specifically, the first trading day (FTD) of each month versus the monthly expiry day of Futures and Options (EXP). The research establishes, through mathematical validation, that SIP returns are independent of the invested amount. Leveraging historical index data, the study finds that SIPs timed on F&O expiry days consistently outperform FTD-SIPs by 0.5–2.5% annually over short- to medium-term horizons (1–3 years), highlighting a tactical advantage due to expiry-related market volatility. However, this advantage diminishes over longer durations (10–20 years), reaffirming that sustained, long-term investing is the key driver of wealth accumulation. Additionally, the analysis challenges prevailing industry narratives by revealing that the 20-year compounded annual growth rate (CAGR) of Nifty 50 SIPs stands at approximately 6.7% (pre-tax), significantly lower than the widely advertised 12–15% returns. This research offers novel insights for retail investors and financial planners on the role of investment timing in SIP strategies within the Indian equity market.
@@ I Introduction
Systematic Investment Plans (SIPs) have emerged as a cornerstone of retail investment in India’s equity markets, offering a structured and disciplined path to long-term wealth creation. Designed to reduce the emotional and cognitive burden of investment decisions, SIPs enable individuals to invest fixed amounts at regular intervals, thereby benefiting from rupee-cost averaging and compounding returns over time. This approach is particularly appealing in India–a rapidly growing economy with a rising middle class seeking accessible wealth-creation tools [36, 23]. Among the available investment avenues, the Nifty 50 Index has emerged as a widely adopted benchmark for systematic investment plans (SIPs) in India. The index comprises 50 large-cap companies across 13 sectors and is considered a proxy for the Indian equity market and the broader economy [19]. According to recent data from the Association of Mutual Funds in India (AMFI), as of February 2025, more than 100 million investors participate in SIPs. Among those, 4.456 million are new SIP accounts, further contributing approximately INR 25,999 crore monthly to mutual funds [2]. This marks a significant rise from early 2021, underscoring the growing popularity of the SIP route among retail investors [3]. While SIPs are lauded for their simplicity and disciplined approach to investing, the timing of these investments remains a relatively underexplored aspect. It is commonly observed that many SIPs are scheduled for execution on the first trading day of each month, a practice likely influenced by salary credit cycles and the perception that early-month investments are financially prudent. However, this uniformity in investment timing may inadvertently expose investors to suboptimal market entry points, potentially affected by institutional order flows and other market dynamics.
Futures and Options (F&O) expiry days—typically the last Thursday of each month—are characterized by heightened trading activity and market turbulence. Empirical studies indicate that expiry-day volumes surge significantly, with price pressures emerging around the event. Vipul (2005) finds that volumes increase abnormally and prices tend to dip just before expiry, followed by sharp gains the day after, largely due to arbitrage activity and constraints in the cash market [32]. These short-term dislocations may offer systematic entry opportunities, particularly for SIPs timed around expiry. While past research–such as the study by [31]–confirms the long-run outperformance of SIPs over lump-sum strategies, these studies do not evaluate the tactical potential of aligning SIP execution with F&O expiry cycles. This gap persists in both academic and practitioner literature regarding whether systematic volatility exposure via expiry-day investments can enhance returns without increasing investor burden.
This study addresses these gaps through a 22-year comparative analysis (2003–2024) of SIP performance under two execution schedules: (i) FTD-SIP (first trading day investments), and (ii) EXP-SIP (F&O expiry day investments). Using the Nifty 50 index as the benchmark, we assess return differentials across horizons ranging from short-term (1–3 years) to long-term (20 years). Our findings reveal three critical insights:
@@
    Timing matters–EXP-SIPs outperform FTD-SIPs by 0.5–2.5% annually in shorter horizons, demonstrating tactical advantages from expiry-day volatility
@@
    Performance is timing-sensitive–Market behavior at entry points dominates the impact of absolute investment amounts
@@
    Myths vs. reality–The widely advertised 12-15% historical returns in index fund [13, 30, 11], contrast sharply with empirical 20-year CAGR of 6.7% (pre-tax), exposing discrepancies in industry narratives.
While expiry-day SIPs offer measurable advantages over short horizons, this benefit fades beyond 10–12 years—underscoring that long-term, consistent participation remains the dominant driver of wealth accumulation. The remainder of this paper is structured as follows: Section 2 reviews theoretical and empirical literature on SIP performance and timing. Section 3 outlines the methodology applied to 22 years of Nifty 50 index data. Section 4 presents theoretical and empirical results comparative performance results across different investment durations. Section 5 discusses broader implications and limitations of the findings, and Section 6 concludes the study.

### v2

@@ Abstract
Systematic Investment Plans (SIPs) are a primary vehicle for retail equity participation in India, yet the impact of their intra-month timing remains underexplored. This study offers a 22-year (2003–2024) comparative analysis of SIP performance in the Nifty 50 index, contrasting the conventional first-trading-day (FTD-SIP) strategy with an alternative aligned to monthly Futures and Options expiry days (EXP-SIP). Using a multi-layered statistical framework—including non-parametric tests, effect size metrics, and stochastic dominance—we uncover two key findings. First, EXP-SIPs outperform FTD-SIPs by 0.5–2.5% annually over short-to-medium-term horizons (1–5 years), with Second-Order Stochastic Dominance (SSD) confirming the EXP-SIP as the preferred choice for all risk-averse investors. Second, we establish boundary conditions for this advantage, showing it decays and becomes economically negligible over longer horizons (10–20 years), where compounding and participation dominate. Additionally, the study challenges the prevalent “12% return” narrative in Indian equity markets, finding that the 20-year pre-tax CAGR for Nifty 50 SIPs is closer to 6.7%. These findings carry significant implications for investor welfare, financial product design, and transparency in return reporting.
@@ I Introduction
Systematic Investment Plans (SIPs) have emerged as a cornerstone of retail investment in India’s equity markets, offering a structured and disciplined path to long-term wealth creation. Designed to reduce the emotional and cognitive burden of investment decisions, SIPs enable individuals to invest fixed amounts at regular intervals, thereby benefiting from rupee-cost averaging and compounding returns over time. This approach is particularly appealing in India–a rapidly growing economy with a rising middle class seeking accessible wealth-creation tools [46, 33]. Among the available investment avenues, the Nifty 50 Index has emerged as a widely adopted benchmark for systematic investment plans (SIPs) in India. The index comprises 50 large-cap companies across 13 sectors and is considered a proxy for the Indian equity market and the broader economy [29]. According to recent data from the Association of Mutual Funds in India (AMFI), as of February 2025, more than 100 million investors participate in SIPs. Among those, 4.456 million are new SIP accounts, further contributing approximately INR 25,999 crore monthly to mutual funds [2]. This marks a significant rise from early 2021, underscoring the growing popularity of the SIP route among retail investors [3]. Despite their popularity, a critical dimension of SIPs remains underexamined: the intra-month timing of execution. A vast number of SIPs are scheduled by default for the first trading day of the month, a convention likely driven by salary credit cycles. However, this uniformity may inadvertently expose a majority of investors to systematically unfavorable market entry points if predictable microstructure effects are present.
One such predictable pattern occurs around monthly Futures and Options (F&O) expiry days—typically the last Thursday of each month—which are marked by elevated trading volumes and short-term price dislocations. Prior research documents significant volume spikes and systematic price dips just before expiry, often followed by sharp rebounds driven by institutional arbitrage and liquidity constraints in the cash market [41]. This raises a compelling question: can aligning SIP execution with this predictable expiry-day volatility yield superior outcomes? While earlier studies confirm the long-run superiority of SIPs over lump-sum strategies [40], they do not explore the tactical potential of intra-month timing. This gap persists in both academic and practitioner discourse. Consequently, whether expiry-related volatility can be harnessed as a source of alpha for systematic investors remains an open empirical question.
This paper addresses this gap through a 22-year comparative analysis (2003–2024) of two SIP timing strategies using the Nifty 50 index: (i) the FTD-SIP, executed on the first trading day of the month, and (ii) the EXP-SIP, executed on the monthly F&O expiry day. We evaluate relative performance across a range of horizons and uncover three key findings:
@@
    A Robust Timing Advantage Exists: The EXP-SIP strategy generates a statistically and economically significant outperformance over the FTD-SIP strategy in 1- to 5-year horizons, proving that intra-month timing is not irrelevant.
@@
    The Advantage is Horizon-Dependent: The nature of the outperformance is dynamic. We demonstrate that the EXP-SIP strategy achieves Second-Order Stochastic Dominance, making it preferable for any risk-averse investor, but the drivers of this advantage shift as the investment horizon lengthens.
@@
    Long-Term Returns are Widely Misrepresented: Our data directly challenges the widely marketed 12–15% return narrative. The actual 20-year pre-tax CAGR for the Nifty 50 is closer to 6.7%, revealing a critical disconnect between industry claims and empirical reality.
Notably, the tactical advantage of the EXP-SIP strategy diminishes over multi-decade horizons, reinforcing the foundational principle that the discipline of consistent, uninterrupted participation remains the dominant driver of long-term wealth. The remainder of this paper is structured as follows: Section 2 reviews theoretical and empirical literature on SIP performance and timing. Section 3 outlines the methodology applied to 22 years of Nifty 50 index data. Section 4 presents theoretical and empirical results comparative performance results across different investment durations. Section 5 discusses broader implications and limitations of the findings, and Section 6 concludes the study.